OpenStudy (anonymous):

(5x^2 - 3x + 1) plus (x^2 - 7) = minus (6x + 2) = plus (4x^2 - x + 8) = ?

OpenStudy (anonymous):

if i remember right that is a compound equation

OpenStudy (anonymous):

no wait a minute why do you put an = sign?

OpenStudy (anonymous):

I've gotten so far: (5x^2 - 3x + 1) + (x^2 - 7) 5x^2(x^2 - 7) = 5x^4 - 35x^2 3x(x^2 - 7) = 3x^3 - 21x (5x^4 - 35x^2) + (3x^3 - 21x) = ?

OpenStudy (anonymous):

It tells us to follow the trails, performing the operations as indicated.

OpenStudy (anonymous):

and why at the end you got a ? And what trails?

OpenStudy (anonymous):

I used the question mark to mean I'm not sure what do do next, or what does the problem equal. The trails are each part of the problem. i.e. (5x^2 - 3x + 1) plus (x^2 - 7), like that, then you move on to the next part of the problem, like after you get the answer to (5x^2 - 3x + 1) plus (x^2 - 7) you go ahead and minus it by (6x + 2).

OpenStudy (anonymous):

Oh isee :D

OpenStudy (anonymous):

Yeah. :)

OpenStudy (anonymous):

(5x^2 - 3x + 1) + (x^2 - 7) = ----------------------- 6x^2 -3x +1 all i did was add like terms :) + -6x - 2) = notice here i took -(6x + 2) and distribute the - ----------------------- 6x^2-9x -1 here i add like terms + (4x^2 - x + 8) = ----------------------------- 10x^2-10x+7

OpenStudy (anonymous):

Thank you so much!